to the nearest tenth of a degree, find the sizes of the acute angles in the right triangle whose hypotenuse is 2.5 times as long as its short leg.: to the nearest tenth of a degree, find the sizes of the acute angles in the right triangle whose hypotenuse is 2.5 times as long as its short leg.
Answer by vleith(1238)

Assume the short side is adjacent to the hypotenuse.
cos(z) = adjacent/hypotenuse = 1/2.5
To find z, use arccosine(1/2.5) = z = 23.6
So one angle is 23.6 degrees and the other one is 180-90-23.6 = 66.4degrees
Question 148271: to the nearest tenth of a degree, find the sizes of the acute angles in the right triangle whose hypotenuse is 2.5 times as long as its short leg.: to the nearest tenth of a degree, find the sizes of the acute angles in the right triangle whose hypotenuse is 2.5 times as long as its short leg.
Answer by Nate(3495)

drawing( 400, 300, -10, 10, -10, 10, triangle( -8, -4, 8, -4, 8, 0 ), locate( 0, -0.5, 2.5x ), locate( 8.5, -1.5, x), locate( -4, -3, O), locate( 7, -0.5, Q ) )
Let's begin with angle O. We need a trig identity that would include the opposite leg and the hypotenuse. _SOH_ CAH TOA
sin O = x / 2.5x
sin O = 1 / 2.5
sin O = 0.4
O = 23.6 degrees
Angle Q = 90 - Angle O
Angle Q = 90 - 23.6
Angle Q = 66.4 degrees